# “Niñas Atómicas” (Atomic Girls): An initiative that generates opportunities for young girls in STEM

### v1

@@
We report on an initiative that seeks to encourage high school girls to develop critical thinking and transferable skills widely used in scientific work, as well as to generate a concrete space of opportunities for girls to experience how real science is done. Our “Niñas Atómicas” workshop combines the teaching of particle physics, electronics, programming and scientific methodology through building and operating a dedicated experiment: a muon counter. Girls from all over Chile can apply to this workshop, where every year they are guided by female scientists for two weeks. We report on the contents and methodology of our workshop and provide details on how to build the muon detector counter experiment. We report results on muon flux and proper lifetime, two muon properties which can be extracted from the data collected by the girls with the muon detectors they built themselves. Insights into the girl’s experiences during the 2024 and 2025 editions of the workshop are also detailed, with the aim to contribute to the wider physics education research and outreach communities.
@@ I Introduction
Particularly in Chile, gender scoring gaps at the high-school level in mathematics Radovic2018; PerezMejias2021 may correlate with gender stereotypes DelRio2016, lack of motivation, confidence or limited access to resources and educational opportunities TIMMS2019. In this context, our science workshop “Niñas Atómicas” (Atomic Girls) Atomicas aims to provide a concrete space of opportunities for high-school girls to experience how real science is done [^1] They are thought by scientists, they built their own experiment, take data with it, analyze it, come up with a scientific question, and then report on their research results. More details in section III.. Our objectives are twofold: i) foster critical thinking and transferable scientific skills useful in our modern world, such as programming or a basic understanding of electronic systems, and ii) that this experience aids in their decision-making process by giving them a concrete understanding of what is science and how science is done, before choosing a major [^2] Having completed STEM related courses at the high-school level helps define a “STEM-readiness” ChoiceSTEM metric, which can impact the chances of entering a STEM program..
Our workshop revolves around an experiment that counts muons (the “muon counter”). In section II we motivate and describe the experiment. In section III, we describe the contents and methodology of the workshop. In section IV we present results on muon properties based on the data the girls took after building their detectors, as well as results addressing the perception the girls had after finishing the workshop. We conclude in section V hoping that the ideas our initiative brings may be replicated or inspire others alike.
@@ II The muon experiment
Experimental setup: The muon counter includes two plastic scintillators, each equipped with a light-sensing silicon photomultiplier (SiPM). The scintillators are wrapped in a reflective foil to concentrate the light emission to the SiPM. The detector also includes two hexagonal casings and a main case, made from plastic rolls (filament) of PLA plastic printed with a standard 3D printer. It is important that the filament is black, as other colors can sometimes let light seep through. The print quality should also be good to make the piece as dense as possible, or alternatively, the infill parameters can be increased so that the piece acts as a black box [^3] With printers that provide a high-quality finish, this adjustment is often unnecessary.. This is so the SiPMs are covered in a dark, non-reflective surface in order to reduce external light contamination to the muon signal. A scintillating material has the property of (re-)emitting energy in the form of photons when a charged particle (as a muon) deposits energy in it. These photons are then processed by the SiPMs, triggering a Geiger discharge and creating a measurable current. The current then goes through a custom-made [^4] Owned and developed by the Center for Theoretical and Experimental Particle Physics (CTEPP) at Universidad Andrés Bello. printed circuit board (PCB), which is the physical platform that tailors and processes the signal, and where the electronic components are installed. The signal from the PCB then enters an Arduino UNO from where we can record the data to an SD card. The Arduino is then connected to a computer, and through the Arduino IDE Software Arduino, a count of the detected muons will be recorded, together with a time stamp at which each muon was detected. In our muon counter design, we require a coincidence between two detector channels to declare a valid muon event. Each detector’s analog signal is passed through a discriminator (comparator) to produce a clean logic pulse. These pulses are inputs to a fast AND gate, which outputs a logic “high” only if both inputs go high within a short time window (on the order of nanoseconds). This coincidence requirement suppresses noise and random single-detector triggers: only when the two SiMPs fire nearly simultaneously is a muon detection registered. The data is saved to a .txt file for further analysis. The components of the muon experiment are shown in Fig. 1.
Figure 1: Components of our muon counter detector experiment. The scintillators in d) and e) are wrapped in aluminum foil and are to be placed on top of the SiPMs in f) and g) when the detector is assembled. The main detector case is shown in a) and its hexagonal casings in b) and c) will contain the SiPMs when closing the detector. The PCB is shown in h), which will be connected to the Arduino in i). The Arduino will be connected to the SD card reader in j).
@@ III The workshop
Our “Niñas Atómicas” workshop has been held every Chilean winter since 2022, lasting two weeks. Girls from all regions of the country can apply online. The criteria to participate was to complete an online application [^5] The online application form was advertised in social media, and was embedded within the Qualtrics Software Qualtrics. They were asked to provide their socio-demographics information, to answer questions on their scientific motivations as well as technical questions intended to measure their scientific skills and critical thinking. A separate article is currently in preparation based on the application instrument we created (see Ref. EstudioImpacto). as well as being between 15 and 16 years of age. Selection was random among all girls that satisfied the criteria. This article presents the results from 50 participants selected in 2024 (21 selected from 70 applicants) and 2025 (29 selected from 257 applicants). The selection process and workshop methodology implemented was the same for the 2024 and 2025 editions.
@@
We detail in what follows the specific sessions of the workshop and what we teach. Throughout all sessions, in selected moments, we ask open questions for the girls to answer in order to diagnose their previous knowledge and to understand their level of comprehension of each session. They can also ask questions anytime.
@@
In the second session we explain what a muon is, how it is created and what are its main properties. We start by emphasizing that not all particles live inside atoms, and tell them how muons are created in cosmic ray air showers and how they were discovered by Carl Anderson y Seth Neddermeyer PhysRev.51.884 . We then explain the properties of the muon: its large mass compared to the electron, its instability via the weak force, its proper lifetime and its large lab frame decay distance, allowing us to detect them.
@@
The third and final electronics session is devoted to explain the characterization of the path a muon will undertake as it travels through the muon counter detector. The different detector components in Figure 1 are explained highlighting their relevance in making the experiment work. The different connections of the SiMPs to the PCB are detailed, as well as how the signal is processed. When there is a coincidence of two nearly simultaneous pulses, the PCB counts a signal. This means a muon passed through both scintillators (placed at each of the detector casings, one at the top and the other one at the bottom, see Fig. 1), which is when a muon is detected. Furthermore, the Arduino is explained detailing how we can save the data with its SD card reader module. The electrical connections detailed in Figure 2 are explained, so they know beforehand what they will need to connect before assembling the detectors themselves. We finish by explaining the role of potentiometers to avoid failures in the PCB and to make sure we feed the detector with the adequate current and voltage.
@@
Figure 2: Diagram showing the electrical connections of our muon counter detector experiment.
@@
Hands-on detector assembly: The girls (together with one parent or guardian) travel to the university in Santiago to work for two days in the detector assembly and data taking. They work in groups of 4 or 5, led by a tutor when assembling the experiment, which takes approximately three to four hours. We provide all components from Figure 1 as well as gloves, PTFE tape, lubricant, screws, cables and a multimeter per group. An oscilloscope is available in the lab as well, so girls can see muons right after they start taking data with their detectors.
The detector assembly starts with the plastic scintillators. The girls wear gloves to first wrap them in aluminum foil. They then place the scintillators on top of the SiPMs. When doing this, the uncovered side should be touching the sensor, and the aluminum covered side should point to the centers of the main case when placing the SiPM inside the hexagonal casings. A small amount of lubricant is applied to the central square of each SiPM when placed inside the hexagonal casings (see also Figure 1).
They then connect the SiPMs to the PCB, where two 3-connector female-to-female cables are needed. Two pins correspond to the power supply voltage and the other one corresponds to the measured signal. When connecting the cables that link the scintillators to the rest of the circuit, the opening of the detector cases should be aligned with the SiPMs cable entries. Screws can be used to secure the cover of the casings after placing the cables. It is important to keep track of which cable corresponds to ground “GND”, signal “S” and positive “++” on the SiPMs before closing the detector, as these need to match the “GND”, “Signal” and “H_Voltage” entries on the PCB as labeled in Figure 2. The pins “GND” and “S” are labeled on the right and left of the SiPMs in f) and g) in Figure 1. On the PCB in h), the ground is the unlabeled entry right below the labels “SiPM 1” and “SiPM 2”.
The next step is to connect the PCB to the Arduino, for which three female-to-male cables are needed. Two of them correspond to the power supply voltage and the other one corresponds to the measured signal (with “GND”, “3,3V” and “Signal” labels in Figure 2). The signal, after being processed by the PCB, comes out through the port labeled “TTL out” (this is written on top of the PCB, see Figure 1). Using a female-to-male cable, this output is connected to the digital input D3 labeled $\sim$3” on the Arduino in Figure 2.
The following step is to supply power to the circuit by connecting the Arduino to a computer with a USB cable or directly to a power socket with a transformer. The girls then prepare the multimeter and calibrate the voltage of the potentiometers of the PCB, which correspond to the three blue boxes labeled (1), (2) and (3) on the PCB in Figure 1. When measuring voltages with the multimeter, one of the probes is always touching a ground point on the PCB (“GND” label in the lower right corner of the PBC) and the other to the point on the PCB where the voltage is to be measured. These measurement points correspond to the circles labeled “TCom1” and “TCom2” right below potentiometers (1) and (2), and towards the left of the “29.4 V” label below (3). For potentiometers (1) and (2) the voltage increases when turning the screw clockwise, and should be set to 2.18 V. For potentiometer (3) the voltage increases when turning the screw counterclockwise and should be set to 29.4 V.
Moving forward, six more cables are needed to connect the SD card reader to the Arduino, with “GND-GND”, “VCC-5V”, “MISO-12”, “MOSI-$\sim$11”, “SCK-13” and “CS-$\sim$10” connections, detailed in Figure 2.
@@
The full assembly of the experiment is shown in Figure 3. A “step-by-step” tutorial (in Spanish) that aids the girls in the detector assembly can be found in AtomicasGithub. After assembly, the girls install the Arduino IDE Software in their computers, following the instructions in Ref Arduino. The program registers an index for each muon together with the time (in seconds) each muon was detected, which the girls can see in real time. A .txt file will be saved to the SD card with this data. We detail a specific format for them to name their files specifying their data taking conditions to avoid confusions. The girls can also take the detectors outside the lab and plug it into the mains of a different place using a transformer to take data under different conditions (i.e. another room or at the top of a building inside campus), after monitoring with their computers that the detectors are working properly. If so, a count of around one muon per minute (at sea level) is expected.
@@
Figure 3: Full assembly of our muon counter detector experiment. The detector components and electrical connections are shown in Figure 1 and 2, respectively.
@@
Programming: Four online sessions of about one hour each are dedicated to explain how they can analyze the data they took with their muon counter experiments. These are distributed in two two-hour modules during the first two days of the second week of the workshop. After each modulus, they have hands-on programming sessions in the afternoon (in breakout rooms over zoom), rounding up our two days devoted to programming.
@@
The lessons start with the importance of using programming tools as an agile, clear and streamlined way to communicate with computers and to perform dedicated tasks as analyzing data. We emphasize that learning to program today is as essential as learning a language like English, since it broadens access to knowledge and allows students to draw their own conclusions about natural phenomena, such as counting muons. We provide everyday examples they use as social media, that make use of algorithms written in the language of computers. We also clearly state that the main objective is to plot their different data sets taken under different conditions (i.e. muon fluxes when placing their detectors at different heights or other different conditions they are curious about).
@@
Scientific Methodology: Two sessions lasting about one hour (the first in person and the second one online) are devoted to explaining how science is done. The first lecture is given on the second in-person day of the workshop, after the girls assembled and tested their muon detectors (so they can start reflecting early on what to do with their data and under which conditions they will take it). We start the first session with a real story on an anomalous measurement in the time of flight of neutrinos being produced at CERN and then detected by the OPERA experiment in Gran Sasso [^8] See Ref. Opera for consecutive CERN press releases on the story. See Ref. OPERA:2011ijq for the OPERA measurement (the first preprint on arXiv evidences the anomaly). Also, this particular story on how science works (and others), can also be found in Tim Lewens book in Ref. lewens2015meaning.. OPERA detected in 2011 an early arrival time of neutrinos when compared to the speed of light in vacuum OPERA:2011ijq. One interpretation of this measurement became popular at that time, that neutrinos could travel faster than the speed of light, breaking the principles of Einstein’s special relativity if it were to be true [^9] We explained earlier in the particle physics sessions what a neutrino is and the importance of special relativity in the measurement of muons.. Throughout this story, we emphasize three key messages: i) the importance of peer review and validation by the community ii) that the ability to confirm or falsify a result (or even a statement) is necessary if we want to know the truth about something, and that this process requires multiple tests and hard work (sometimes of hundreds of people), as, paraphrasing Carl Sagan, “extraordinary claims require extraordinary evidence” and that “only through inquire we can discover truth” Sagan1979 iii) the importance of predictive power in a scientific theory, and the need of agreement of these predictions with experiment, considering all experiments to date, no “matter how beautiful your theory is or how smart you are”, quoting Richard Feynman Feynman1964 without invalidating previous knowledge and especially if you don’t have enough evidence). In this later message we stress on the value of making predictions or formulating hypothesis, and that this is a brave endeavor. We emphasize that, in the words of Tim Lewens, “what makes something a genuine piece of science is its potential vulnerability to refutation” lewens2015meaning. We exemplify predictions of both the Standard Model of Particle Physics and General Relativity (i.e. particle masses, how frequently particles are produced by the LHC, the bending of light and gravitational waves), two scientific theories that have passed the tests of time.
@@
Along our second session, we explain how in science we document the process of doing it. We show them examples of real scientific papers, and encourage them to ask themselves a question related to the muons they measured. One of the most common and straightforward question they can ask is : “Does the muon count changes if I place the detector at different heights?” . Another common question is “Does the muon count changes if I put some type of material (it could be a rock, or a magnet) on top of the detector?”. The girls then report their findings to such questions individually. They make a scientific report in the following two afternoons after the programming sessions (we also devote the second-to-last day of the workshop for them to meet online with their groups to answer questions on their reports). They can do their reports in any format, including a presentation, an infographic or a more standard written report, as long as they write it in a conservative and objective language and include the expected contents: i) a clear question or hypothesis they put to the test ii) a brief introduction that provides context and motivates their question iii) an analysis section where they demonstrate what they did and how, including plots or tables with the data they took iv) a summary section, where they must link their findings to their concrete scientific question or hypothesis. They must comment on weather their curiosity can - or can not - be answered with their data. Or to which extent, particularly, if their experiment failed. We encourage them to think on what other possible measurements they could do in order to investigate further to provide answers to their questions, stressing that an important aspect of scientific work is to embrace doubts, and so they can evidence that their work indeed opens more questions.
@@
Close: The last day of the workshop starts with dedicated feedback, where each tutor explains what each of the girls in their group did well according to everything we taught them, as well as where there is room for improvement. The main objective of this session is that the girls receive comments of their scientific reports, in a constructive way and in a safe space, as this is done for each group were only their tutors and teammates are present. After this, and as a last session of the workshop, we invite an “Atomic Woman” to give the girls a talk on their work and scientific career. The goal is that the girls can listen to the motivations and unique journey of a female scientist from a different field. Our invited speakers from the 2024 and 2025 workshops gave talks on Data Analysis and Machine Learning, and on Quantum Optics. The workshop ends with an informal “Q&\&A” and farewell session.

### v2

@@
We report on an initiative that seeks to encourage high school girls to develop critical thinking and transferable skills widely used in scientific work, as well as to generate a concrete space of opportunities for girls to experience how real science is done. Our “Niñas Atómicas” workshop combines the teaching of particle physics, electronics, programming and scientific methodology through building and operating a dedicated experiment: a muon counter. Girls from all over Chile can apply to this workshop, where every year they are guided by female scientists for two weeks. We report on the contents and methodology of our workshop and provide details on how to build the muon detector. We report results on muon flux and proper lifetime, two muon properties which can be extracted from the data collected by the girls with the muon detectors they built themselves. Insights into the girl’s experiences during the 2024 and 2025 editions of the workshop are also detailed, with the aim to contribute to the wider physics education research and outreach communities.
@@ I Introduction
Particularly in Chile, gender scoring gaps at the high-school level in mathematics Radovic2018; PerezMejias2021 may correlate with gender stereotypes DelRio2016, lack of motivation, confidence or limited access to resources and educational opportunities TIMMS2019. In this context, our science workshop “Niñas Atómicas” (Atomic Girls) Atomicas aims to provide a concrete space of opportunities for high-school girls to experience how real science is done [^1] They are taught by scientists, they built their own experiment, take data with it, analyze it, come up with a scientific question, and then report on their research results. More details in section III.. Our objectives are twofold: i) foster critical thinking and transferable scientific skills useful in our modern world, such as programming or a basic understanding of electronic systems, and ii) that this experience aids in their decision-making process by giving them a concrete understanding of what is science and how science is done, before choosing a major [^2] Having completed STEM related courses at the high-school level helps define a “STEM-readiness” ChoiceSTEM metric, which can impact the chances of entering a STEM program..
Our workshop revolves around a hands-on experiment that counts muons using a dedicated muon detector. In section II we motivate and describe the experiment. In section III, we describe the contents and methodology of the workshop. In section IV we present results on muon properties based on the data the girls took after building their detectors, as well as results addressing the perception the girls had after finishing the workshop. We conclude in section V hoping that the ideas our initiative brings may be replicated or inspire others alike.
@@ II The muon experiment
Experimental setup: The muon detector we use in our workshop includes two plastic scintillators, each equipped with a light-sensing silicon photomultiplier (SiPM). The scintillators are wrapped in a reflective foil to concentrate the light emission to the SiPM. The detector also includes two hexagonal casings and a main case, made from plastic rolls (filament) of PLA plastic printed with a standard 3D printer. It is important that the filament is black, as other colors can sometimes let light seep through. The print quality should also be good to make the piece as dense as possible, or alternatively, the infill parameters can be increased so that the piece acts as a black box [^3] With printers that provide a high-quality finish, this adjustment is often unnecessary.. This is so the SiPMs are covered in a dark, non-reflective surface in order to reduce external light contamination to the muon signal. A scintillating material has the property of (re-)emitting energy in the form of photons when a charged particle (as a muon) deposits energy in it. These photons are then processed by the SiPMs, triggering a Geiger discharge and creating a measurable current. The current then goes through a custom-made [^4] Owned and developed by the Center for Theoretical and Experimental Particle Physics (CTEPP) at Universidad Andrés Bello. printed circuit board (PCB), which is the physical platform that tailors and processes the signal, and where the electronic components are installed. The signal from the PCB then enters an Arduino UNO from where we can record the data to an SD card. The Arduino is then connected to a computer, and through the Arduino IDE Software Arduino, a count of the detected muons will be recorded, together with a time stamp at which each muon was detected. In our muon detector design, we require a coincidence between two detector channels to declare a valid muon event. Each detector’s analog signal is passed through a discriminator (comparator) to produce a clean logic pulse. These pulses are inputs to a fast AND gate, which outputs a logic “high” only if both inputs go high within a short time window (on the order of nanoseconds). This coincidence requirement suppresses noise and random single-detector triggers: only when the two SiMPs fire nearly simultaneously is a muon detection registered. The data is saved to a .txt file for further analysis. The components of the muon experiment are shown in Fig. 1.
Figure 1: Components of our muon detector. The scintillators in d) and e) are wrapped in aluminum foil and are to be placed on top of the SiPMs in f) and g) when the detector is assembled. The main detector case is shown in a) and its hexagonal casings in b) and c) will contain the SiPMs when closing the detector. The PCB is shown in h), which will be connected to the Arduino in i). The Arduino will be connected to the SD card reader in j). Figure adapted from a photograph taken by Ignacio García.
@@ III The workshop
Our “Niñas Atómicas” workshop has been held every Chilean winter since 2022, lasting two weeks. Girls from all regions of the country can apply online. This article presents the results from 50 participants selected in 2024 (21 selected from 70 applicants) and 2025 (29 selected from 257 applicants). The selection process and workshop methodology implemented was the same for the 2024 and 2025 editions. The criteria to participate was to complete an online application [^5] The online application form was advertised in social media, and was embedded within the Qualtrics Software Qualtrics. They were asked to provide their socio-demographics information, to answer questions on their scientific motivations as well as technical questions intended to measure their scientific skills and critical thinking. A separate article is currently in preparation based on the application instrument we created (see Ref. EstudioImpacto). as well as being between 15 and 16 years of age. Selection was random among all girls that satisfied the criteria.
@@
In what follows, we detail the specific sessions of the workshop and what we teach. Throughout all sessions, in selected moments, we ask open questions for the girls to answer in order to diagnose their previous knowledge and to understand their level of comprehension of each session. They can also ask questions anytime.
@@
In the second session we explain what a muon is, how it is created and what are its main properties. We start by emphasizing that not all particles live inside atoms, and tell them how muons are created in cosmic ray air showers and how they were discovered by Carl Anderson and Seth Neddermeyer PhysRev.51.884 . We then explain the properties of the muon: its large mass compared to the electron, its instability via the weak force, its proper lifetime and its large lab frame decay distance, allowing us to detect them.
@@
The third and final electronics session is devoted to explain the characterization of the path a muon will undertake as it travels through the muon detector. The different detector components in Figure 1 are explained highlighting their relevance in making the experiment work. The different connections of the SiMPs to the PCB are detailed, as well as how the signal is processed. When there is a coincidence of two nearly simultaneous pulses, the PCB counts a signal. This means a muon passed through both scintillators, which is when a muon is detected. Furthermore, the Arduino is explained detailing how we can save the data with its SD card reader module. The electrical connections detailed in Figure 2 are explained, so they know beforehand what they will need to connect before assembling the detectors themselves. We finish by explaining the role of potentiometers to avoid failures in the PCB and to make sure we feed the detector with the adequate current and voltage.
@@
Figure 2: Diagram showing the electrical connections of our muon detector. Detailed instructions can be found in AtomicasGithub.
@@
Hands-on detector assembly: The girls (together with one parent or guardian) travel to the university in Santiago to work for two days in the detector assembly and data taking. They work in groups of 4 or 5, led by a tutor when assembling the experiment, which takes approximately three to four hours. We provide all components from Figure 1 as well as gloves, PTFE tape, lubricant, screws, cables and a multimeter per group. An oscilloscope is available in the lab as well, so girls can evidence a muon signal right after they start taking data with their detectors.
@@
The detector assembly starts by placing the plastic scintillators and the SiPM inside the detector’s hexagonal casings. The students then electrically connect the SiPMs to the PCB, and the PCB to the Arduino. They then supply power to the circuit by connecting the Arduino to a computer with a USB. The electrical connections are shown in Figure 2. The full assembly of the experiment is shown in Figure 3. A “step-by-step” tutorial (in Spanish) made by our tutors, as well as a version in English, can be found in AtomicasGithub. After assembly, the girls install the Arduino IDE Software in their computers, following the instructions in Ref Arduino. The program registers an index for each muon together with the time (in seconds) at which each muon was detected, which the girls can see in real time. A .txt file will be saved to the SD card with this data. We detail a specific format for them to name their files specifying their data taking conditions to avoid confusions. When monitoring with a computer that the detectors are working properly, a count of around one muon per minute (at sea level) is expected. We note that the detectors could also be used outside the lab without a computer, by plugging them directly into the mains of a different place using a transformer.
@@
Figure 3: Full assembly of our muon detector. The detector components and electrical connections are shown in Figure 1 and 2, respectively. We thank Ignacio García for taking this photograph.
@@
Programming: Four online sessions of about one hour each are dedicated to explain how they can analyze the data they took with their muon detector. These are distributed in two two-hour modules during the first two days of the second week of the workshop. After each modulus, they have hands-on programming sessions in the afternoon (in breakout rooms over Zoom), rounding up our two days devoted to programming.
@@
The lessons start with the importance of using programming tools as an agile, clear and streamlined way to communicate with computers and to perform dedicated tasks as analyzing data. We emphasize that learning to program today is as essential as learning a language like English, since it broadens access to knowledge and allows students to draw their own conclusions about natural phenomena, such as counting muons. We provide everyday examples they use as social media, that make use of algorithms written in the language of computers. We also clearly state that the main objective is to plot their different data sets taken under different conditions (e.g. muon fluxes when placing their detectors at different heights or other different conditions they are curious about).
@@
Scientific Methodology: Two sessions lasting about one hour (the first in person and the second one online) are devoted to explaining how science is done. The first lecture is given on the second in-person day of the workshop, after the girls assembled and tested their muon detectors (so they can start reflecting early on what to do with their data and under which conditions they will take it). We start the first session with a real story on an anomalous measurement in the time of flight of neutrinos being produced at CERN and then detected by the OPERA experiment in Gran Sasso [^8] See Ref. Opera for consecutive CERN press releases on the story. See Ref. OPERA:2011ijq for the OPERA measurement (the first preprint on arXiv evidences the anomaly). Also, this particular story on how science works (and others), can also be found in Tim Lewens book in Ref. lewens2015meaning.. OPERA detected in 2011 an early arrival time of neutrinos when compared to the speed of light in vacuum OPERA:2011ijq. One interpretation of this measurement became popular at that time, that neutrinos could travel faster than the speed of light, breaking the principles of Einstein’s special relativity if it were to be true [^9] We explained earlier in the particle physics sessions what a neutrino is and the importance of special relativity in the measurement of muons.. Throughout this story, we emphasize three key messages: i) the importance of peer review and validation by the community ii) that the ability to confirm or falsify a result (or even a statement) is necessary if we want to know the truth about something, and that this process requires multiple tests and hard work (sometimes of hundreds of people), as, paraphrasing Carl Sagan, “extraordinary claims require extraordinary evidence” and that “only through inquire we can discover truth” Sagan1979 iii) the importance of predictive power in a scientific theory, and the need of agreement of these predictions with experiment, considering all experiments to date, no “matter how beautiful your theory is or how smart you are”, quoting Richard Feynman Feynman1964 without invalidating previous knowledge and especially if you don’t have enough evidence). In this later message we stress on the value of making predictions or formulating hypothesis, and that this is a brave endeavor. We emphasize that, in the words of Tim Lewens, “what makes something a genuine piece of science is its potential vulnerability to refutation” lewens2015meaning. We exemplify predictions of both the Standard Model of Particle Physics and General Relativity (e.g. particle masses, how frequently particles are produced by the LHC, the bending of light and gravitational waves), two scientific theories that have passed the tests of time.
@@
Along our second session, we explain how in science we document the process of doing it. We show them examples of real scientific papers, and encourage them to ask themselves a question related to the muons they measured. One of the most common and straightforward question they can ask is : “Does the muon counts change if I place the detector at different heights?” . Another common question is “Does the muon counts change if I put some type of material (it could be a rock, or a magnet) on top of the detector?”. The girls then report their findings to such questions individually. They make a scientific report in the following two afternoons after the programming sessions (we also devote the second-to-last day of the workshop for them to meet online with their groups to answer questions on their reports). They can do their reports in any format, including a presentation, an infographic, or a more standard written report, as long as they write it in a conservative and objective language and include the expected contents: i) a clear question or hypothesis they put to the test ii) a brief introduction that provides context and motivates their question iii) an analysis section where they demonstrate what they did and how, including plots or tables with the data they took iv) a summary section, where they must link their findings to their concrete scientific question or hypothesis. They must comment on weather their curiosity can - or can not - be answered with their data. Or to which extent, particularly, if their experiment failed. We encourage them to think on what other possible measurements they could do in order to investigate further to provide answers to their questions, stressing that an important aspect of scientific work is to embrace doubts, and so they can evidence that their work indeed opens more questions.
@@
Closure: The last day of the workshop starts with dedicated feedback, where each tutor explains what each of the girls in their group did well according to everything we taught them, as well as where there is room for improvement. The main objective of this session is that the girls receive comments of their scientific reports, in a constructive way and in a safe space, as this is done for each group where only their tutors and teammates are present. After this, and as a last session of the workshop, we invite a Chilean scientist or an “Atomic Woman” to give the girls a talk on their work and scientific career. The goal is that the girls can listen to the motivations and unique journey of a female scientist from a different field. Our invited speakers from the 2024 and 2025 workshops gave talks on Data Analysis and Machine Learning, and on Quantum Optics. The workshop ends with an informal “Q&\&A” and farewell session.